### v1

@@ 3.4 Results from FRG
We first elucidate the fixed point structure of the FNS equation within an even simpler approximation, which consists in considering only the (non-perturbative) flow of the non-linear dimensionless coupling parameter $\bar{g}_{\kappa}=\kappa^{d-4}\chi\lambda/\nu_{\kappa}^{2}$, with $\nu_{\kappa}\equiv\bar{\nu}_{\kappa}(0,0)$. Since we are interested in the inviscid fixed point, corresponding to $\bar{g}_{\kappa}\to\infty$, we define $\hat{w}_{\kappa}=\bar{g}_{\kappa}/(1+\bar{g}_{\kappa})$. Its flow is given by
$\partial_{s}\hat{w}_{\kappa}=\hat{w}_{\kappa}(1-\hat{w}_{\kappa})\big{(}d-2+2\eta_{\kappa}\big{)}\,,\quad\hbox{with}\;\;\eta_{\kappa}=-\partial_{s}\ln\nu_{\kappa}\,.$ (27)
@@
This equation can be straightforwardly integrated. The result in shown in Fig. 1, where it is compared with the equivalent flow for the 1D Burgers equation from Ref. [23]. From any initial condition $\hat{w}_{\Lambda}$ at $\kappa=\Lambda$ ($s=\ln(\kappa/\Lambda)=0$), the flow reaches, as $\kappa\to 0$ ($s\to-\infty$), an IR fixed point, with $\hat{w}^{\rm FNS}_{*}=0$ which is the attractive FNS fixed point. To evidence UV fixed points, one can run the flow backwards, from a very small initial $\kappa$ (very large negative initial $s$) to $\kappa\to\Lambda$ ($s\to 0$), such that the stability of the fixed points are exchanged. The backwards flow always reaches in the UV a new fixed point, with $\hat{w}^{\rm IFNS}_{*}=1$, which corresponds to the IFNS fixed point, as illustrated on Fig. 1(b). This demonstrates the existence of this fixed point. In contrast, for the 1D Burgers flow, while the flow always leads in the IR to a non-Gaussian fixed point with a finite $10<\hat{w}^{\rm KPZ}_{*}<$ 1 which is the KPZ fixed point, it can reach in the UV two different UV fixed points depending on the initial value of the coupling: the IB fixed point for $\hat{w}_{0}>\hat{w}^{\rm KPZ}_{*}$, and the EW one for $\hat{w}_{0}<\hat{w}^{\rm KPZ}_{*}$, as illustrated on Fig. 1(a).
@@
Figure 1: FRG flows of the non-linear coupling $\hat{w}_{\kappa}$ for the stochastic (a) 1D Burgers and (b) 3D Navier-Stokes equations: in red flows from $s=\ln(\kappa/\Lambda)=0$ to $s\to-\infty$ ($\kappa\to 0$) towards the IR fixed points for different initial conditions $\hat{w}_{\Lambda}$; in blue flows from $s\gtrsim-20$ to $s\to 0$ ($\kappa\to\Lambda$) towards the UV fixed points for different initial conditions $\hat{w}_{0}$.
@@
This unveils the whole fixed point structure of model A, which is sketched in Fig. 2, again in correspondence with the one of the Burgers–KPZ equation. The FNS fixed point describes the large distance and time, i.e. the IR modes (symbolised by the red-shaded area in the figure), of the viscous fluid at rest. The IFNS fixed point describes the inviscid fluid at equilibrium, but it also controls the large (UV) modes of a viscous fluid, the range of these modes (symbolised by the blue-shaded area in the figure) extending as the viscosity decreases.
@@
Figure 2: Sketch of the fixed point structure of the stochastic 1D Burgers and 3D Navier-Stokes equations in the presence of thermal noise. The red points represent attractive (IR) fixed points and the blue points repulsive (UV) ones. The red-shaded (resp. blue-shaded) area hence symbolises the IR (resp. UV) modes controlled by the IR (resp. UV) fixed points.
@@
In order to obtain more quantitative results, and in particular the correlation function, we numerically solve the complete set of flow equations (23), (26), using two complementary grids [52]. A first grid is used for the small momentum sector, where the flow equation (26) is integrated using dimensionless variables with a similar numerical scheme as described in Refs. [23, 49]. The large momentum flow equation (23) is integrated on a second grid using dimensionful variables, with the input at each scale $\kappa$ of $J_{\kappa}(\omega)$ computed on the dimensionless grid. The full numerical procedure is detailed in Ref. [56]. This scheme allows us to compute the correlation function $\bar{C}(\omega,\mathbf{p})$ on the whole momentum range. To evidence scaling regimes, we extract from it the half decay frequency defined such that $\bar{C}(\omega_{1/2}(q),\mathbf{p})=\bar{C}(0,\mathbf{p})/2$. The result is displayed on Fig. 3. Two markedly different scaling regimes are observable, thus unveiling the existence of two underlying fixed points. At small momenta, one finds $\omega_{1/2}\sim q^{z}$ with $z=2$ the FNS diffusive exponent. This region is hence controlled by the IR FNS fixed point as expected. For large viscosities (eg $\nu=1$), only this scaling is visible. For large momenta, when the viscosity decreases, one finds another scaling regime where $\omega_{1/2}\sim q^{z}$ with $z=1$. This regime is all the more extended than the viscosity is small.
@@
Figure 3: Half decay frequency for different initial values of the viscosity. The IR modes show a diffusive FNS scaling $\omega_{1/2}\sim p^{2}$. The IFNS scaling, characterised by $\omega_{1/2}\sim p$, appears for UV modes when the viscosity decreases.
@@ 4.1 Numerical methods
The stochastically forced Navier-Stokes equations (1)– (3) are numerically solved using $2/3$ de-aliased Fourier spectral methods [57]. We thus use the discrete Fourier transform $\mathbf{v}(t,\mathbf{x})=\sum\hat{\mathbf{v}}(t,\mathbf{k})e^{i\mathbf{k}\cdot\mathbf{x}}$ of a 3D spatially periodic velocity field. Performing a Galerkin truncation on (1)–(3), i.e. ${\hat{\mathbf{v}}}(t,\mathbf{k})=0$ for $k_{\alpha}k_{\alpha}\geq k_{\rm max}^{2}$ (where $k_{\rm max}$ is $1/3$ of the resolution), generates the following finite system of ordinary differential equations for the complex variables ${\hat{\mathbf{v}}}(t,\mathbf{k})$ ($\mathbf{k}$ being a 3D vector of relative integers $(k_{1},k_{2},k_{3})$)
@@
At large times, the truncated 3D Euler equations is known to reach an absolute equilibrium, which is a statistically stationary exact Gaussian solution, with energy spectrum $E(k)=ck^{2}$ [12]. When $\nu$ and $D$ are non-zero, identical statistically stationary solutions also exist. A realisation of such a solution, with $u_{rms}=\sqrt{2E}=\frac{1}{2}$, is used as initial data to start our numerical simulations. When $\nu$ is non-zero, $D$ is adjusted so that the corresponding stationary solution also has $u_{rms}=\frac{1}{2}$ (see [17] for details on a similar setting for the 1D Burgers equation).
We use fields with Taylor-Green symmetries (see Ref. [13] for details). Time-stepping of the deterministic terms of (28) is performed using a $4^{\rm th}$-order Runge-Kutta method and the random forcing term is added independently. Two-dimensional slices of the 3D computed Fourier-space velocity field are output periodically every $\delta t$ to files. Post-processing the files, time-correlation functions and spatio-temporal power spectra up to the Nyquist pulsation $\omega_{\rm max}=\pi/\delta t$ are estimated using standard algorithms [58]. The runs presented here use spatial resolution of $256^{3}$ and $u_{rms}=\frac{1}{2}$.The inviscid run uses the output frequency $\delta t=0.1$ and a total integration time of $t=1024$. The viscous run with $\nu=0.016$ uses $\delta t=0.0125$ and a total integration time of $t=128$.
@@ 4.2 Scalings of correlation times
The red circles  [^2] Red circles standing right on top of the $0.5$ contour line validates the interpolation scheme used to obtain the logarithmic representation from the raw numerical data that is equally spaced both in time and wavenumber. indicate the points $\tau_{1/2}(k)$ obtained independently by solving for each value of $k$ the implicit definition $\bar{C}(\tau_{1/2}(k),k)={\frac{1}{2}}\bar{C}(0,k)$. The left panel of Fig. 4 corresponds to the inviscid case $\nu=D=0$. The black dashed lines indicate the $t\sim k^{-1}$ and $t\sim k^{-2}$ scaling laws, which are both observable. In fact, the small-time behaviour of the correlation scaling with timescale $t\sim k^{-1}$ was already predicted in the inviscid limit in Ref. [12, 13], in agreement with the FRG exact result for small time delays (24). This scaling is apparent on the level lines from 0.9 to 0.5 which closely follow this law. Interestingly, at larger times, another scaling emerges, indicated by the level line $0.1$ which is definitely steeper and nearer to a $t\sim k^{-2}$ behaviour. This is an evidence for the FRG prediction for the large time delays behaviour (24). On the right panel of Fig. 4, the forced Navier-Stokes equation obtained at $\nu=0.016$ is displayed. The black dashed line indicates the theoretical scaling $t\sim k^{-2}$, corresponding to the diffusive FNS behaviour, which is well-followed for this large viscosity.
@@
The scaling behaviour can also be observed on the spectrum $S(k,\omega)$, defined as the Fourier transform in time of $\bar{C}(t,k)/\bar{C}(0,k)$. The numerical noise on this quantity can be reduced by considering $\omega S(k,\omega)$ instead [57, 58]. Fig. 5 shows the contour plot of $\omega S(k,\omega)$. Scaling laws again correspond to straight lines in the $(\log(\omega),\log(k))$ plane. The left panel of Fig. 5 displays the inviscid ($\nu=0$) scalings, where the black dashed lines indicate $\omega\sim k$ and $\omega\sim k^{2}$ scaling laws. It is apparent that both scalings are present: $\omega\sim k^{2}$ for small $\omega$ and $\omega\sim k$ for large $\omega$. On the right panel, the viscous scaling obtained at $\nu=0.016$ is displayed. The black dashed line indicate the expected theoretical $\omega\sim k^{2}$ diffusive law, which is again well followed.
@@
One can study as well the emergence of the $z=1$ scaling varying the viscosity. As highlighted in Fig. 3 obtained with FRG, one expects that at finite but non-zero viscosity, this scaling also appears at large wavenumbers, on a range all the more extended that the viscosity is small. We show on the left panel of Fig. 6 the half decay time $\tau_{1/2}(k)$ obtained from numerical simulations from $\nu=0.016$ to $\nu=0$, and on the right panel the same data compensated by $k$, i.e $k\tau_{1/2}(k)$, such that the $z=1$ scaling appears in this representation as a plateau. Note that at the highest momenta, all the curves saturate due to truncation effects. For the largest viscosities, only the diffusive regime $z=2$ is observed, while as the viscosity decreases, the behaviour of the large momentum mode bends towards a $z=1$ scaling. The plateau is visible on the right panel for viscosities $\nu\lesssim 0.0022$. The numerical results are thus in perfect accordance with the FRG prediction.
@@
Figure 6: Left panel: half decay time $\tau_{1/2}$ of the correlations versus wavenumber $k$ for different viscosities obtained from numerical simulations. At the largest viscosity, only the viscous scaling $\tau_{1/2}\sim k^{-2}$ is visible, indicated by the solid line. At the smallest viscosities, the behaviour at large wavenumbers shows the inviscid scaling $\tau_{1/2}\sim k^{-1}$, indicated by the dashed line. Right panel: same as in left panel, but compensated by $k$, i.e $k\,\tau_{1/2}(k)$ versus $k$. Markers correspond on both panels to: $\nu=0.016$: $o$, $\nu=0.0044$: asterisk, $\nu=0.0022$: $+$, $\nu=0.0011$: diamond, $\nu=0.00055$: x and $\nu=0$: squares (colors from red to blue).

### v2

@@ 3.4 Results from FRG
We first elucidate the fixed point structure of the FNS equation within an even simpler approximation, which consists in considering only the (non-perturbative) flow of the non-linear dimensionless coupling parameter $\bar{g}_{\kappa}=\kappa^{d-2}\chi\lambda^{2}/\nu_{\kappa}^{2}$, with $\nu_{\kappa}\equiv\bar{\nu}_{\kappa}(0,0)$. Since we are interested in the inviscid fixed point, corresponding to $\bar{g}_{\kappa}\to\infty$, we define $\hat{w}_{\kappa}=\bar{g}_{\kappa}/(1+\bar{g}_{\kappa})$. Its flow is given by
$\partial_{s}\hat{w}_{\kappa}=\hat{w}_{\kappa}(1-\hat{w}_{\kappa})\big(d-2+2\eta_{\kappa}\big)\,,\quad\hbox{with}\;\;\eta_{\kappa}=-\partial_{s}\ln\nu_{\kappa}\,$, (27)
@@
and the flow $\partial_{s}\nu_{\kappa}$ is obtained from (26) evaluated at zero external frequency $\omega=0$ and momentum $\mathbf{p}=0$. This equation can be straightforwardly integrated numerically. The result in shown in Fig. 1, where it is compared with the equivalent flow for the stochastic 1D Burgers equation from Ref. [28]. From any initial condition $\hat{w}_{\Lambda}$ at $\kappa=\Lambda$ ($s=\ln(\kappa/\Lambda)=0$), the flow reaches, as $\kappa\to 0$ ($s\to$-$\infty$), an IR fixed point corresponding to $\hat{w}^{\rm FNS}_{*}=0$. This is the attractive FNS fixed point. To evidence UV fixed points, one can run the flow backwards, starting from a very small initial $\kappa\equiv\kappa_{0}$ (very large negative initial $s$) to $\kappa\to\Lambda$ ($s\to 0$), such that the stability of the fixed points are exchanged. As illustrated on Fig. 1(b), the backwards flow always reaches in the UV a new fixed point, corresponding to $\hat{w}^{\rm IFNS}_{*}=1$, which we call IFNS. This demonstrates the existence of this fixed point. In contrast, for the 1D stochastic Burgers equation, the RG flow always leads in the IR to a non-Gaussian fixed point with a finite $0<\hat{w}^{\rm KPZ}_{*}<1$ which is the KPZ fixed point. Running the flow backwards from a very large negative value of $s$ (very small $\kappa_{0}$) to $s\to 0$, the flow can reach two different UV fixed points depending on the initial value of the coupling: the IB fixed point for $\hat{w}_{\kappa_{0}}>\hat{w}^{\rm KPZ}_{*}$, and the EW one for $\hat{w}_{\kappa_{0}}<\hat{w}^{\rm KPZ}_{*}$, as illustrated on Fig. 1(a).
@@
Figure 1: FRG flows of the non-linear coupling $\hat{w}_{\kappa}$ for (a) the stochastic 1D Burgers and (b) the stochastic 3D Navier-Stokes equations: in red flows from $s=\ln(\kappa/\Lambda)=0$ to $s\to$-$\infty$ ($\kappa\to 0$) towards the IR fixed points for different initial conditions $\hat{w}_{\Lambda}$; in blue flows from small initial $\kappa_{0}$ (large negative $s\lesssim-20$) to $s\to 0$ ($\kappa\to\Lambda$) towards the UV fixed points for different initial conditions $\hat{w}_{\kappa_{0}}$.
@@
This unveils the whole fixed point structure of model A, which is sketched in Fig. 2, again in correspondence with the one of the 1D Burgers–KPZ equation. The FNS fixed point describes the behaviour at large distance and time, i.e. the behaviour of the IR modes (symbolised by the red-shaded area in the figure), of the viscous fluid at rest. The IFNS fixed point describes the inviscid fluid at equilibrium, but it also controls the large (UV) modes of a viscous fluid, the range of these modes (symbolised by the blue-shaded area in the figure) extending as the viscosity decreases.
@@
Figure 2: Sketch of the fixed point structure of the stochastic 1D Burgers and stochastic 3D Navier-Stokes equations in the presence of thermal noise. The red points represent attractive (IR) fixed points and the blue points repulsive (UV) ones. The red-shaded (resp. blue-shaded) area hence symbolises the IR (resp. UV) modes controlled by the IR (resp. UV) fixed points.
@@
In order to obtain more quantitative results, and in particular the correlation function, we numerically solve the complete set of flow equations (23), (26), using two complementary grids [57]. A first grid is used for the small momentum sector, where the flow equation (26) is integrated using dimensionless variables with a similar numerical scheme as described in Refs. [28, 54]. The large momentum flow equation (23) is integrated on a second grid using dimensionful variables, with the input at each scale $\kappa$ of $J_{\kappa}(\omega)$ computed on the dimensionless grid. The full numerical procedure is detailed in Ref. [62]. This scheme allows us to compute the correlation function $\bar{C}(\omega,\mathbf{p})$ on the whole momentum range. To evidence scaling regimes, we extract from it the half decay frequency defined such that $\bar{C}(\omega_{1/2}(p),\mathbf{p})=\bar{C}(0,\mathbf{p})/2$. The result is displayed on Fig. 3. Two markedly different scaling regimes are observable, thus unveiling the existence of the two underlying fixed points. At small momenta, one finds $\omega_{1/2}\sim q^{z}$ with $z=2$ the FNS diffusive exponent. This region is hence controlled by the IR FNS fixed point as expected. For large viscosities (eg $\nu=1$), only this scaling is visible. For large momenta, when the viscosity decreases, one finds another scaling regime where $\omega_{1/2}\sim q^{z}$ with $z=1$. This regime is all the more pronounced than the viscosity is small.
@@
Figure 3: Half decay frequency for different initial values of the viscosity. The IR modes show a diffusive FNS scaling $\omega_{1/2}\sim p^{2}$. The IFNS scaling, characterised by $\omega_{1/2}\sim p$, appears for UV modes when the viscosity decreases. In this figure, the momentum $p$ is measured in units of $\Lambda$ and the viscosity $\nu$ in units of $(\Lambda\chi)^{1/2}$.
@@ 4.1 Numerical methods
The stochastic Navier-Stokes equation with thermal noise (1)–(3) are numerically solved using $2/3$ de-aliased Fourier spectral methods [63]. We thus use the discrete Fourier transform $\mathbf{v}(t,\mathbf{x})=\sum\hat{\mathbf{v}}(t,\mathbf{k})e^{i\mathbf{k}\cdot\mathbf{x}}$ of a 3D spatially periodic velocity field. Performing a Galerkin truncation on (1)–(3), i.e. ${\hat{\mathbf{v}}}(t,\mathbf{k})=0$ for $k_{\alpha}k_{\alpha}\geq k_{\rm max}^{2}$ (where $k_{\rm max}$ is $1/3$ of the resolution), generates the following finite system of ordinary differential equations for the complex variables ${\hat{\mathbf{v}}}(t,\mathbf{k})$ ($\mathbf{k}$ being a 3D vector of relative integers $(k_{1},k_{2},k_{3})$)
@@
At large times, the truncated 3D Euler equations is known to reach an equilibrium, which is a statistically stationary exact Gaussian solution, with energy spectrum $E(k)=ck^{2}$ [11]. When $\nu$ and $D$ are non-zero, identical statistically stationary solutions also exist. A realisation of such a solution, with $u_{rms}=\sqrt{2E}=\frac{1}{2}$, is used as initial data to start our numerical simulations. When $\nu$ is non-zero, $D$ is adjusted so that the corresponding stationary solution also has $u_{rms}=\frac{1}{2}$ (see [20] for details on a similar setting for the 1D Burgers equation).
We use fields with Taylor-Green symmetries (see Ref. [13] for details). Time-stepping of the deterministic terms of (28) is performed using a $4^{\rm th}$-order Runge-Kutta method and the random thermal noise term is added independently. Two-dimensional slices of the 3D computed Fourier-space velocity field are output periodically every $\delta t$ to files. Post-processing the files, time-correlation functions and spatio-temporal power spectra up to the Nyquist pulsation $\omega_{\rm max}=\pi/\delta t$ are estimated using standard algorithms [64]. The runs presented here use spatial resolution of $256^{3}$ and $u_{rms}=\frac{1}{2}$.The inviscid run uses the output frequency $\delta t=0.1$ and a total integration time of $t=1024$. The viscous run with $\nu=0.016$ uses $\delta t=0.0125$ and a total integration time of $t=128$.
@@ 4.2 Scalings of correlation times
The red circles  [^3] Red circles standing right on top of the $0.5$ contour line validates the interpolation scheme used to obtain the logarithmic representation from the raw numerical data that is equally spaced both in time and wavenumber. indicate the points $\tau_{1/2}(k)$ obtained independently by solving for each value of $k$ the implicit definition $\bar{C}(\tau_{1/2}(k),k)={\frac{1}{2}}\bar{C}(0,k)$. The left panel of Fig. 4 corresponds to the inviscid case $\nu=D=0$. The black dashed lines indicate the $t\sim k^{-1}$ and $t\sim k^{-2}$ scaling laws, which are both observable. In fact, the small-time scaling behaviour of the correlation with timescale $t\sim k^{-1}$ was already predicted in the inviscid limit in Ref. [11, 13], in agreement with the FRG exact result for small time delays (24). This scaling is apparent on the level lines from 0.9 to 0.5 which closely follow this law. Interestingly, at larger times, another scaling emerges, indicated by the level line $0.1$ which is definitely steeper and nearer to a $t\sim k^{-2}$ behaviour. This is an evidence for the FRG prediction for the large time-delay behaviour (24). On the right panel of Fig. 4, the results for the stochastic Navier-Stokes equation obtained at $\nu=0.016$ is displayed. The black dashed line indicates the theoretical scaling $t\sim k^{-2}$, corresponding to the diffusive FNS behaviour, which is well-followed for this large viscosity.
@@
The scaling behaviour can also be observed on the spectrum $S(k,\omega)$, defined as the Fourier transform in time of $\bar{C}(t,k)/\bar{C}(0,k)$. The numerical noise on this quantity can be reduced by considering $\omega S(k,\omega)$ instead [63, 64]. Fig. 5 shows the contour plot of $\omega S(k,\omega)$. Scaling laws again correspond to straight lines in the $(\log(\omega),\log(k))$ plane. The left panel of Fig. 5 displays the inviscid scalings, where the black dashed lines indicate $\omega\sim k$ and $\omega\sim k^{2}$ scaling laws. It is apparent that both scalings are present: $\omega\sim k^{2}$ for small $\omega$ and $\omega\sim k$ for large $\omega$. On the right panel, the viscous scaling obtained at $\nu=0.016$ is displayed. The black dashed line indicates the expected theoretical $\omega\sim k^{2}$ diffusive law, which is again well followed.
@@
One can study as well the emergence of the $z=1$ scaling varying the viscosity. As highlighted in Fig. 3 obtained with FRG, one expects that at finite but non-zero viscosity, this scaling also appears at large wavenumbers, all the more pronounced that the viscosity is small. We show on the left panel of Fig. 6 the half decay time $\tau_{1/2}(k)$ obtained from numerical simulations from $\nu=0.016$ to $\nu=0$, and on the right panel the same data compensated by $k$, i.e $k\tau_{1/2}(k)$, such that the $z=1$ scaling appears in this representation as a plateau. Note that at the highest momenta, all the curves saturate due to truncation effects. For the largest viscosity, only the diffusive regime $z=2$ is observed, while as the viscosity decreases, the behaviour of the large momentum mode bends towards a $z=1$ scaling. The plateau is visible on the right panel for viscosities $\nu\lesssim 0.0022$. The numerical results are thus in perfect accordance with the FRG prediction.
@@
Figure 6: Left panel: half decay time $\tau_{1/2}$ of the correlations versus wavenumber $k$ for different viscosities obtained from numerical simulations. At the largest viscosity, only the viscous scaling $\tau\sim k^{-2}$ is visible, indicated by the solid line. At the smallest viscosities, the behaviour at large wavenumbers shows the inviscid scaling $\tau\sim k^{-1}$, indicated by the dashed line. Right panel: same as in left panel, but compensated by $k$, i.e $k\,\tau_{1/2}(k)$ versus $k$. Markers correspond on both panels to: $\nu=0.016$: $o$, $\nu=0.0044$: asterisk, $\nu=0.0022$: ++, $\nu=0.0011$: diamond, $\nu=0.00055$: x and $\nu=0$: squares (colors from red to blue).
5 Relevance of the IFNS scaling in physical systems
An important question is whether the unveiled scaling behaviour is observable in a real fluid at equilibrium. In Fig. 3, the momentum $p$ is measured in units of $\Lambda$ and the viscosity $\nu$ in units of $(\Lambda\chi)^{1/2}$. The scale $\Lambda^{-1}$ represents the length scale at which the hydrodynamical description is valid, it should be chosen large enough with respect to the typical microscopic length, for instance $\Lambda^{-1}\sim 10\lambda­_{\text{mfp}}$, where $\lambda­_{\text{mfp}}$ is the mean free path. This would set for air the units to be $1.5\times 10^{4}\text{cm}^{-1}$ for momentum and 0.0007 cm2/s for viscosity, and for water $4\times 10^{6}\text{cm}^{-1}$ for momentum and 0.0004 cm2/s for viscosity at T=300T=300K. To observe the $z=1$ scaling, we need to set $\bar{g}_{\Lambda}\gg 1$, or $\nu\ll 1\times 10^{-4}\text{cm}^{2}$/s, and this behaviour emerges for wavenumbers starting from $10^{4}\text{cm}^{-1}$ for air and $10^{5}\text{cm}^{-1}$ for water. However, even without substituting specific constants, the estimation of the viscosity for these fluids (eg. in Ref. [14], Eq. (2.26), which agrees with numerical calculation of viscosity at the scale of theămean free path in Ref. [26]) shows that $\bar{g}_{\Lambda}$ remains small for molecular fluids as long as the scale $\Lambda^{-1}$ is larger than the mean free path. Thus, the $z=1$ scaling probably does not lie within the physical region for molecular fluids at equilibrium. On the other hand, it becomes observable when the fluid is turbulent.